A machine with a cost of $52,000 has an estimated residual value of $4,246 and an estimated life of 5 years or 16,907 hours. What is the amount of depreciation for the second full year, using the double-declining-balance method?
Answer #1

Straight-line rate = (100/5) = 20%

Double declining rate = (2*20)% = 40%

Depreciation for first year = 52,000 * 40% = 20,800

Depreciation for second year = (52,000-20,800) * 40%

= 12,480